OpenStudy (anonymous):

A baseball player slides into third base with an initial speed of 5.0 m/s. If the coefficient of kinetic friction between the player and the ground is 0.45, how far does the player slide before coming to rest?

OpenStudy (theeric):

Hi! Have you attempted a solution, @Bing2009 ?

OpenStudy (anonymous):

The mass of the player makes no difference in this problem because more mass means more normal force down and more horizontal force from momentum, so they cancel each other d = V^2 / 2 u *g d = distance in meters to stop V = velocity of player u = coefficient of kinetic friction = 0.45 g = gravity acceleration = 9.81 m/s^2 d = (5 m/s)^2 / 2 * 0.45 * 9.81 m/s^2 d = 2.8 meters the player slides to a stop

OpenStudy (theeric):

I agree! The trick is deriving that equation. Let us know if you have any questions, @Bing2009